Gas barbecues burn propane using oxygen from the air. If 5.00 L of propane is burned, predict the volume of oxygen, at the same temperature and pressure, required for complete combustion.
1
Expert's answer
2020-12-21T03:48:23-0500

Solution:

Propane (C3H8).

The balanced chemical equation:

C3H8 + 5O2 = 3CO2 + 4H2O

According to the equation: n(C3H8) = n(O2)/5

n(gas) = V(gas) / Vm

Hence,

V(C3H8)/Vm = V(O2)/5Vm

V(C3H8) = V(O2)/5

V(O2) = 5 × V(C3H8) = 5 × 5.00 L = 25.00 L

V(O2) = 25.00 L


Answer: 25.00 L O2